# DeVAIC: A Tool for Security Assessment of AI-generated Code

### v1

@@ 5.1 NL Prompts Details
We designed these prompts starting from those provided by three publicly available datasets about code generation, i.e. SecurityEval SecurityEval_prompts , LLMSecEval LLMSecEval_prompts and CodeXGLUE CodeXGLUE_prompts . First of all, we modified these prompts by avoiding the insertion of implementation details. This line of action allows the model to generate code according to their discretion without any influence on how to implement a task. Moreover, we adopted the perspective of an ordinary, nonmalicious user who requested the model to generate code for a given task, which means that our prompts do not contain explicit triggers to induce the prediction of unsafe code.
@@
Write a Python code with a get_name() function that gets the name parameter from the URL by using request module by importing flask and opens the file names/name returning a file object
@@
Table 3 displays two examples of NL prompts and a sample of vulnerable code generated by the models we employed. Row #1 showcases a model’s prediction that uses the request.args.get() function of the Flask module. The only validation produced in the code (i.e., the only one defined by the user in the prompt) is to verify if the output parameter of the request.args.get() function is an empty variable. Nevertheless, it is not sufficient because, even if this parameter is not empty, it may contain dangerous characters (i.e., “../”, “../..”, etc.) that can allow a malicious user to access files or directories outside of the intended directory (i.e., name/). The CWE associated with this kind of vulnerability is CWE-022, commonly known as “Path Traversal”. This CWE is related to the Broken Access Control category of OWASP’s Top 10. To enhance the security, the code has to ensure that the file path is actually within the expected directory before opening the file itself owasp_pathtrav . We also notice that the generated code snippet lacks completeness, making the application of some static analysis tools not feasible.
Finally, Row #2 shows that the request to evaluate a user input involved the generation of code containing the eval() function. This function is considered dangerous due to the potential injection of malicious code and is related to CWE-095, also known as “Eval Injection”, which belongs to the Injection OWASP category. In fact, the official ast documentation recommends using literal_eval(), a more secure alternative to eval(), that can evaluate only a restricted set of expressions ast_doc .
@@ 5.2 Manual Analysis
As manual classification can be susceptible to errors, three researchers of our research group, who have skills in Python code vulnerability detection, conducted this evaluation independently. Subsequently, they compared their evaluation and performed an in-depth analysis of the few discrepancy cases. This collaborative effort resulted in a complete alignment, achieving a $100\%$ consensus in the final evaluation.
On average, the four models produced $54\%$ of vulnerable code ($271$ vulnerable code over $500$ predictions), which allows us to comprehensively assess the tool’s detection ability in both vulnerable and non-vulnerable instances, providing a robust and fair measure of their overall performance across various scenarios.
@@ 5.3 Experimental Results
Precision
DeVAIC
0.97
1.00
0.95
0.95
0.97
Bandit
0.89
0.81
0.83
0.82
0.84
CodeQL
0.79
0.86
0.83
0.95
0.85
Semgrep
0.87
0.98
0.94
0.85
0.91
PyT
1.00
0.89
1.00
1.00
0.96
Recall
DeVAIC
0.95
0.96
0.90
0.86
0.92
Bandit
0.70
0.69
0.51
0.58
0.62
CodeQL
0.36
0.54
0.42
0.26
0.39
Semgrep
0.55
0.69
0.58
0.51
0.58
PyT
0.11
0.11
0.08
0.04
0.09
$F_{1}$ Score
DeVAIC
0.96
0.98
0.92
0.90
0.94
Bandit
0.78
0.74
0.63
0.68
0.72
CodeQL
0.49
0.67
0.56
0.41
0.54
Semgrep
0.67
0.81
0.72
0.64
0.71
PyT
0.20
0.20
0.16
0.08
0.16
Accuracy
DeVAIC
0.95
0.98
0.93
0.89
0.94
Bandit
0.77
0.74
0.71
0.69
0.73
CodeQL
0.56
0.70
0.68
0.58
0.63
Semgrep
0.69
0.82
0.78
0.67
0.74
PyT
0.48
0.50
0.56
0.46
0.50
To provide context for the evaluation, we compared DeVAIC’s performance with a baseline. In our analysis, we utilized CodeQL version v2.16.4 and the two Security test suites sec_testsuite2 ; sec_testsuite1 of queries for Python. We also used Bandit version 1.7.7, Semgrep version 1.61.1, and python-taint module version 0.42, also known as PyT. We remark again that several snippets generated by the models were incomplete, preventing us from directly analyzing them with some well-known static analysis tools, such as Bandit, CodeQL, or others. At first, we analyzed the model predictions with DeVAIC. Then, we modified the incomplete ones ($65$ in total) by inserting the import statement when lacking to compare DeVAIC’s evaluation results with the state of the art.
Considering the average values for all models, Table 4 shows that DeVAIC achieved high values for each metric. Regarding Precision, DeVAIC reaches $97\%$, which is comparable with the values achieved by Semgrep [^4] When we used Semgrep, we executed a total of 1291 rules from the official ruleset registry for Python semgrep_registry . ($91\%$) and PyT ($96\%$). Furthermore, DeVAIC achieved a maximum Recall of $92\%$ on average, surpassing the baseline with a considerable gap compared to the employed solutions, which obtain a maximum of $62\%$ (i.e., Bandit). Finally, DeVAIC shows superior performance to other state-of-the-art solutions ($\geq 20\%$) with average values of $94\%$ for both $F_{1}$ Score and Accuracy.
We further evaluated if, for each metric (i.e., Precision, Recall, F1 Score, and Accuracy), there is a statistical difference between DeVAIC and the state-of-the-art solutions using the non-parametric Wilcoxon rank sum test. The null hypothesis is that the two samples derive from the same population (i.e., the two populations have equal medians). If the null hypothesis is rejected, the Wilcoxon rank sum test indicates that the two samples are statistically different. As deducible by the considerations explained above, for the Precision, DeVAIC is statistically different only with Bandit while obtaining optimum results for all of the other metrics, for which DeVAIC results are statistically different (and better) with every other result obtained with the state of the art. The previously mentioned considerations highlight the outstanding performance of DeVAIC.
@@ 5.4 Computational Cost
We compared the execution time taken by DeVAIC with the other solutions employed. Figure 4 shows the median execution time of each tool in analyzing $125$ code snippets from the four models. Bandit is the faster ($0.67$ seconds), while CodeQL is the slower ($570$ seconds, which corresponds to approximately $9$ minutes). PyT and Semgrep employed $253.40$ and $123.67$ seconds, respectively ($2$ and $4$ minutes approximately). In this context, while the other tools employed minutes on average to analyze the snippets, DeVAIC remains in the realm of seconds, as Bandit does.
@@ 6 Threats to Validity
Rule Creation: The creation of detection rules based on regular expressions could introduce bias, as these rules are derived from patterns observed within a limited set of vulnerable code snippets. This limitation may result in rules that are either too specific or too general, affecting the tool’s accuracy in real-world scenarios. To mitigate this threat, we employed a diverse dataset of vulnerabilities covering a broad range of CWEs and OWASP categories. We also conducted iterative refinements of our rules by testing them on separate validation sets to ensure they accurately capture the intended patterns without being overly broad or narrow.

### v2

@@ 4.3 Tool Execution
To illustrate how DeVAIC works in practice, we refer to the vulnerable code generated by AI provided in Section 3, in which we have already discussed the vulnerabilities exposed and the issues related to analyzing it using state-of-the-art tools, which is due to the code’s incompleteness. In contrast, the evaluation using DeVAIC highlighted its effectiveness in analyzing incomplete code snippets. Specifically, as previously mentioned, the preliminary step is to convert the code into a single-line format, with the carriage return represented by the “\$\backslash n$” symbol. Subsequently, we ran DeVAIC, obtaining the identification of the OWASP category associated with the vulnerability implemented in the analyzed code, i.e., Software and Data Integrity Failures. Going deeper into the tool functioning, the code triggered a rule that uses the grep command to intercept the vulnerable function yaml.load().
We share DeVAIC and the files to reproduce our experiments on the following URL: https://github.com/dessertlab/DeVAIC.
@@ 5.1 NL Prompts Details
The NL inputs used to query the models are extracted from the test set used in [17]. In this work, the authors queried state-of-the-art code generators to assess whether models generate unsafe code. The authors, which included also security experts, manually constructed the test set by using $100$ NL descriptions of code by combining two benchmark datasets used for evaluating the security of AI-generated code (i.e., SecurityEval [92], and LLMSecEval [103]). Although these $100$ prompts do not contain triggers that force the models to circumvent their security measures for code generation, they still potentially lead the models to generate unsafe code when they are “poisoned”, i.e., fine-tuned with vulnerable corpora. This test set perfectly fits with the scope of our paper since it allows us to evaluate the performance of DeVAIC and other baseline tools on the AI-generated code. To perform a more comprehensive assessment, we enriched the set of NL inputs with an additional set of $25$ examples (for a total of $125$ unique prompts) that require performing operations in the context of Flask (a Python web application framework) and extracted from the CodeXGLUE benchmark [59]. The $125$ prompts cover a comprehensive set of domains, such as GET and POST implementation, file opening, temporary file creation and permissions changing, operations on input read from an external source, etc, without including specific implementation details about how to implement the requested task. In this way, we stimulated the models to generate code snippets representative of real-case scenarios.
@@
Table 3 displays two examples of NL prompts and a sample of vulnerable code generated by the models we employed. Row #1 showcases a model’s prediction that uses the request.args.get() function of the Flask module. The only validation produced in the code (i.e., the only one defined by the user in the prompt) is to verify if the output parameter of the request.args.get() function is an empty variable. Nevertheless, it is not sufficient because, even if this parameter is not empty, it may contain dangerous characters (i.e., “../”, “../..”, etc.) that can allow a malicious user to access files or directories outside of the intended directory (i.e., name/). The CWE associated with this kind of vulnerability is CWE-022, commonly known as “Path Traversal”. This CWE is related to the Broken Access Control category of OWASP’s Top 10. To enhance the security, the code has to ensure that the file path is actually within the expected directory before opening the file itself [76]. We also notice that the generated code snippet lacks completeness, making the application of some static analysis tools not feasible.
Finally, Row #2 shows that the request to evaluate a user input involved the generation of code containing the eval() function. This function is considered dangerous due to the potential injection of malicious code and is related to CWE-095, also known as “Eval Injection”, which belongs to the Injection OWASP category. In fact, the official ast documentation recommends using literal_eval(), a more secure alternative to eval(), that can evaluate only a restricted set of expressions [84].
@@ 5.2 Manual Analysis
Table 4: Occurrences of CWEs in vulnerable snippets generated by each model. The “-” signifies that a specific CWE is not present. CWEs overlapping with Table 1 are in bold.
CWE Google Gemini OpenAI ChatGPT-3.5 Microsoft Copilot GitHub Copilot Total for CWE
CWE-020 13 16 10 10 49
CWE-022 5 5 7 4 21
CWE-078 4 5 2 6 17
CWE-079 3 14 8 4 29
CWE-089 1 - - 2 3
CWE-090 3 - - 1 4
CWE-094 15 14 2 2 33
CWE-095 1 - - - 1
CWE-113 - 1 - - 1
CWE-117 - 2 2 - 4
CWE-209 - 25 - - 25
CWE-215 15 12 2 2 31
CWE-259 6 9 2 2 19
CWE-276 - - 1 - 1
CWE-295 2 - 1 2 5
CWE-319 2 1 2 3 8
CWE-326 3 - 1 1 5
CWE-327 5 3 5 7 20
CWE-330 2 - 1 2 5
CWE-337 2 1 1 - 4
CWE-338 - 1 1 2 4
CWE-347 4 1 - 2 7
CWE-377 - - 1 1 2
CWE-400 5 - - 4 9
CWE-477 - - - 1 1
CWE-489 8 - - - 8
CWE-502 8 8 9 8 33
CWE-601 1 1 1 1 4
CWE-611 5 1 - - 6
CWE-614 - 1 1 - 2
CWE-703 1 1 2 - 4
CWE-732 2 6 3 8 19
CWE-776 1 2 2 - 5
CWE-798 1 6 3 4 14
CWE-918 - - 1 - 1
Total for model 118 136 71 79
As manual classification can be susceptible to errors, it was conducted by a diverse group of $3$ human evaluators, all with a strong background and expertise in cyber security and AI code generators. The group included individuals with varying degrees of professional experience and educational qualifications. In particular, $2$ Ph.D. students and a post-doctoral researcher, all with a computer engineering degree. To minimize the potential for human error, the $3$ human evaluators independently examined each code snippet generated by the four models to check whether it contained or not vulnerabilities, by assigning a score of $1$ or 00, respectively. Then, the evaluators compared their results and performed an in-depth analysis of the few discrepancy cases. The few discrepancies, which consisted of a $\sim 2\%$ on cases, were attributed to human misclassification and subsequently resolved in a complete alignment, achieving a $100\%$ consensus in the final evaluation. Thanks to the diversity and expertise of our evaluators and the iterative process of analysis, we ensured the reliability of our human evaluation process.
On average, the four models produced $54\%$ of vulnerable code ($271$ vulnerable code over $500$ predictions). Furthermore, the same human experts that identified the vulnerabilities in the code generated by models, also checked for the CWE associated with the code. Again, this in-depth analysis was performed independently by consulting the MITRE reports [64]. Afterward, the evaluators compared their results, collaboratively reviewing the few discrepancies encountered, until reaching a 100% consensus. The details about the CWE categories associated with each vulnerable snippet generated by the four models are listed in Table 4. In particular, we marked in bold the $21$ CWEs that overlap with those listed in Table 1, i.e., the CWEs of the samples we used for the rule creation.
The total of $271$ vulnerable snippets generated showed $118$ CWEs for the vulnerable snippets produced by the model Google Gemini, $136$ CWEs for OpenAI ChatGPT-$3.5$, $71$ CWEs for Microsoft Copilot and $79$ CWEs for GitHub Copilot, as shown in Table 4 - “Total for model” row. For each model, there are some vulnerable instances for which the evaluators identified more than a single CWE (i.e., $34\%$ out of vulnerable snippets for Gemini, $53\%$ for ChatGPT, $15\%$ for Microsoft Copilot, and $12\%$ for GitHub Copilot). Some of these CWEs are closely related to each other and allowed us to capture the different characteristics of the vulnerabilities shown in these snippets (e.g., the generated snippet in Table 5 is vulnerable to CWE-$95$ (Eval Injection) and CWE-$209$ (Information Exposure Through an Error Message), the latter is closely related to CWE-$215$ (Insertion of Sensitive Information Into Debugging Code) and CWE-$489$ (Active Debug Code)).
Table 5: Example of CWE labeling for a vulnerable generate snippet.
CWE Vulnerable snippet
[…]
Overall, the $500$ generated code snippets allowed us to comprehensively assess the tool’s detection ability in both vulnerable and non-vulnerable instances, providing a robust and fair measure of their overall performance across various scenarios.
@@ 5.3 Experimental Results
Precision DeVAIC 0.97    1.00    0.95    0.95    0.97
Bandit 0.89    0.81    0.83    0.82    0.84
CodeQL 0.79    0.86    0.83    0.95    0.85
Semgrep 0.87    0.98    0.94    0.85    0.91
PyT 1.00    0.89    1.00    1.00    0.96
ChatGPT-3.5 0.93    0.91    0.77    0.82    0.85
ChaGPT-4 0.71    0.72    0.71    0.68    0.71
Claude-3.5-Sonnet 0.66    0.74    0.74    0.74    0.72
Recall DeVAIC 0.95    0.96    0.90    0.86    0.92
Bandit 0.70    0.69    0.51    0.58    0.62
CodeQL 0.36    0.54    0.42    0.26    0.39
Semgrep 0.55    0.69    0.58    0.51    0.58
PyT 0.11    0.11    0.08    0.04    0.09
ChatGPT-3.5 0.71    0.57    0.68    0.71    0.67
ChatGPT-4 0.77    0.66    0.69    0.72    0.71
Claude-3.5-Sonnet 0.75    0.64    0.78    0.88    0.76
$F_{1}$ Score DeVAIC 0.96    0.98    0.92    0.90    0.94
Bandit 0.78    0.74    0.63    0.68    0.72
CodeQL 0.49    0.67    0.56    0.41    0.54
Semgrep 0.67    0.81    0.72    0.64    0.71
PyT 0.20    0.20    0.16    0.08    0.16
ChatGPT-3.5 0.81    0.70    0.72    0.76    0.75
ChaGPT-4 0.74    0.69    0.70    0.70    0.71
Claude-3.5-Sonnet 0.71    0.69    0.76    0.81    0.74
Accuracy DeVAIC 0.95    0.98    0.93    0.89    0.94
Bandit 0.77    0.74    0.71    0.69    0.73
CodeQL 0.56    0.70    0.68    0.58    0.63
Semgrep 0.69    0.82    0.78    0.67    0.74
PyT 0.48    0.50    0.56    0.46    0.50
ChaGPT-3.5 0.80    0.73    0.75    0.75    0.76
ChatGPT-4 0.68    0.66    0.71    0.65    0.68
Claude-3.5-Sonnet 0.63    0.67    0.76    0.76    0.71
To provide context for the evaluation, we compared DeVAIC’s performance with a baseline. In our analysis, we utilized CodeQL version v2.16.4 and the two Security test suites [22, 23] of queries for Python. We also used Bandit version 1.7.7, Semgrep version 1.61.1, and python-taint module version 0.42, also known as PyT.
As introduced in Section 2, Semgrep uses a pattern-matching approach by executing regular expressions (regex) to search for vulnerable patterns in the code. In the official Registry [96], Semgrep provides ready-to-use configuration files containing regex for vulnerability detection. Instead, tools like CodeQL, Bandit and PyT first model the code under examination with an Abstract Syntax Tree (AST) and then they execute their detection rules on the AST nodes. More in detail, Bandit builds the AST of the code under examination, and runs appropriate plugins (i.e., assert_used, exec_used, set_bad_file_permissions, etc.) [6] against the AST nodes. Once Bandit has finished scanning all the source code, it generates a final report. CodeQL treats code like data, which means that it is necessary to generate a CodeQL database to represent the codebase before running the analysis queries. Even in this case, CodeQL provides public test suites with ready-to-use queries [22, 23]. After the code scanning, CodeQL generates a report. Finally, PyT generates the AST and creates the Control Flow Graph (CFG); then, it passes the CFG to a Framework Adaptor, which will mark the arguments of certain functions (by Flask, Django and other libraries) as tainted sources. In the final step, PyT checks if the output from the tainted sources is sanitized. If not, it raises an alert and generates a report.
We remark again that several snippets generated by the models in our setup were incomplete, so the static analysis tools just mentioned cannot define the AST, thus having issues conducting their detection analyses. At first, we analyzed the model predictions with DeVAIC. Then, we modified the incomplete ones ($65$ in total) by inserting the import statement when lacking to compare DeVAIC’s evaluation results with the state-of-the-art.
Furthermore, we enriched our analysis by comparing DeVAIC even with LLMs-based detection methods, which are nowadays a current trend and a widely used solution in the state of the art to perform vulnerability detection [108, 43]. In particular, we adopted ChatGPT-$3.5$ and ChatGPT-$4$ [70], and the new model Claude-$3.5$-Sonnet [2]. ChatGPT-3.5 and ChatGPT-4 are developed by OpenAI and represent successive advancements in language model performance, with ChatGPT-4 offering enhanced understanding, reasoning, and generation capabilities compared to its predecessors. Claude-3.5-Sonnet, on the other hand, is the new model developed by Anthropic AI, which has set new performance standards in several tasks [3].
We prompted these $3$ models using a Zero-Shot Role-Oriented (ZS-RO) prompt [104], i.e., we assigned to the model the role of vulnerability detection system (i.e., RO), successively asking to perform the vulnerability detection of the $500$ generated snippets with a question about the related CWE (NL prompt: ‘‘You are a vulnerability detection system. Your task is to analyze the following code snippet and identify any potential security vulnerabilities. Specifically, please determine if the code is susceptible to any known Common Weakness Enumerations (CWEs) and provide the corresponding CWE identifier.”
Considering the average values for all models, Table 6 shows that DeVAIC achieved high values for each metric. Regarding Precision, DeVAIC reaches $97\%$, which is comparable with the values achieved by Semgrep [^4] When we used Semgrep, we executed a total of 1291 rules from the official ruleset registry for Python [96]. ($91\%$) and PyT ($96\%$), while surpassing the values achieved by ChatGPT-$4$ ($71\%$) and Claude-$3.5$-Sonnet ($72\%$). Furthermore, DeVAIC achieved a maximum Recall of $92\%$ on average, surpassing the baseline with a considerable gap compared to the employed solutions, which obtain a maximum of $76\%$ (i.e., Claude-$3.5$-Sonnet). Finally, DeVAIC shows superior performance to other state-of-the-art solutions ($\geq 20\%$) with average values of $94\%$ for both $F_{1}$ Score and Accuracy.
Considering the CWEs mapped to the experimental dataset shown in Table 4, we analyzed the snippets detected by DeVAIC and the baseline tools to evaluate how many CWE categories were covered. We found that DeVAIC detected a set of vulnerable snippets associated with $31$ out of the $35$ CWEs listed in Table 4 (i.e., $89\%$), confirming the high performance exhibited by the evaluation metrics in Table 6. Regarding the baseline tools, Bandit detected snippets associated with $18$ CWEs (i.e., $51\%$), while we obtained $22$ CWEs for CodeQL (i.e., $63\%$), $21$ CWEs for Semgrep (i.e., $60\%$), $8$ CWEs for PyT (i.e., $23\%$), $29$ CWEs for Claude-$3.5$-Sonnet (i.e., $83\%$), $19$ CWEs for ChatGPT-$3.5$ (i.e., $54\%$), and $21$ CWEs for ChatGPT-$4$ (i.e., $60\%$). The baseline tools correctly identified some instances of generated snippets labelled with multiple CWEs, justifying the high number of CWEs covered. However, they missed other vulnerable instances, obtaining lower evaluation metrics than DeVAIC, as shown in Table 6.
We further evaluated if, for each metric (i.e., Precision, Recall, F1 Score, and Accuracy), there is a statistical difference between DeVAIC and the state-of-the-art solutions using the non-parametric Wilcoxon rank sum test. The null hypothesis is that the two samples derive from the same population (i.e., the two populations have equal medians). If the null hypothesis is rejected, the Wilcoxon rank sum test indicates that the two samples are statistically different. As deducible by the considerations explained above, for the Precision, DeVAIC is statistically different with Bandit, and with the models ChatGPT-$4$ and Claude-$3.5$-Sonnet. For all of the other metrics, DeVAIC achieves optimal results, which are statistically different (and better) than any other results obtained with the state of the art. The previously mentioned considerations highlight the outstanding performance of DeVAIC.
@@ 5.4 Computational Cost
We compared the execution time taken by DeVAIC with the other solutions employed. Figure 4 shows the median execution time of each tool in analyzing $500$ code snippets from the four models. Bandit is the faster ($0.67$ seconds), while CodeQL is the slower ($570$ seconds, which corresponds to approximately $9$ minutes). PyT and Semgrep employed $253.40$ and $123.67$ seconds, respectively ($2$ and $4$ minutes approximately). In this context, while the other tools employed minutes on average to analyze the snippets, DeVAIC remains in the realm of seconds, as Bandit does.
When using LLM models as vulnerability detectors, the time it takes for analysis depends heavily on human-AI interaction. If we approximate the time for a single detection, using the same prompt but changing snippets from time to time, to be about one minute, and if we analyze $500$ snippets, the total time for a single model would be $500$ minutes ($\sim 8$ hours).
@@ 6 Threats to Validity
Rule Creation: The creation of detection rules based on regular expressions could introduce bias, as these rules are derived from patterns observed within a limited set of vulnerable code snippets. This limitation may result in rules that are either too specific or too general, affecting the tool’s accuracy in real-world scenarios. To mitigate this threat, we employed a diverse dataset of vulnerabilities covering a broad range of CWEs and OWASP categories [95, 80]. We also conducted iterative refinements of our rules by testing them on separate validation sets to ensure they accurately capture the intended patterns without being overly broad or narrow.
Coverage of CWEs: A key threat to the validity of our study concerns the comprehensive coverage of CWEs using detection rules. Given the variability in how a single CWE can manifest across different code patterns, it is challenging to ensure that all possible implementations are adequately covered. This complexity arises because a CWE can present itself in multiple distinct ways, making it difficult to design a definitive set of rules that captures every variation. To mitigate this threat, we implemented 85 detection rules, which is significantly more than the 35 CWEs addressed in our study. This was done to capture a broad spectrum of implementation patterns for each CWE. However, we acknowledge that the possibility remains that some patterns might not be covered by the existing rules, potentially leaving certain vulnerabilities undetected. Despite these challenges, the effectiveness of our detection tool, DeVAIC, is demonstrated by the results obtained from our experimental dataset. DeVAIC successfully detected 91% (248 out of 271) of the vulnerable snippets generated by AI models. This high detection rate suggests that our rule set is robust and capable of identifying a wide range of CWE implementations. However, it is important to note that while this result is promising, it does not guarantee that all potential patterns were captured, as the remaining 9% of vulnerabilities were not detected. Future work will enhance the robustness of our detection tool by continuously updating the rule set to incorporate new CWE patterns, which is crucial for maintaining comprehensive coverage in a dynamic cybersecurity landscape.
@@
Manual Evaluation: A potential threat to the validity of our manual analysis is the subjective nature of human evaluation, which can introduce biases and inconsistencies. Despite the evaluators’ strong backgrounds in cybersecurity and AI code generation, differences in interpretation and judgment could affect the classification of vulnerabilities in the generated code snippets. To address these concerns, we implemented several rigorous measures to mitigate the risk of bias and ensure consistency. First, each code snippet was independently evaluated by three experts (a group comprising two Ph.D. students and a post-doctoral researcher, all with substantial expertise in the relevant domains). This independence in evaluation helps reduce the influence of individual biases, as each evaluator’s assessment is made without knowledge of the others’ judgments. Following the independent evaluations, we conducted a further inspection process for cases with discrepancies, which accounted for only about 2% of the total evaluations. This low discrepancy rate suggests a high level of agreement among the evaluators, indicating robust initial assessments. For the discrepancies that did occur, the evaluators engaged in detailed discussions to reach a consensus, ensuring that any potential misunderstandings or differing interpretations were resolved. This iterative approach helped align the evaluators’ perspectives and provided an opportunity for in-depth analysis, enhancing the overall reliability of the evaluation.